# Towards Robust Real-World Multivariate Time Series Forecasting: A Unified Framework for Dependency, Asynchrony, and Missingness

### v1

@@ Appendix A Implementation Details
A.1 Dataset Configuration under Channel-wise Asynchrony
We conduct long-term forecasting experiments on three modified benchmark datasets adapted to our practical setting, as well as one real-world dataset. In this setting, each channel provides a different number of observations over the same temporal window, reflecting channel-specific sampling periods in the multivariate time series. This configuration more closely mirrors real-world conditions, where sensor deployment strategies and data acquisition policies vary across channels.
ETT1-practical (ETT1)
ETT1-practical is a variant of the commonly used ETTm1 [55] dataset. Unlike the original version, where all channels are sampled uniformly at 15-minute intervals, ETT1-practical introduces channel-specific sampling periods to reflect industrial environments, where sensors operate at different acquisition frequencies depending on their physical properties and monitoring requirements. Specifically, the channels are grouped into two categories: load-related channels and the oil temperature (OT) channel. The load-related channels are downsampled to an hourly rate, representing slowly varying electrical load measurements, while the OT channel retains its original 15-minute resolution to capture high-frequency thermal dynamics.
@@ A.1 Dataset Configuration under Channel-wise Asynchrony
Weather-practical is constructed from the widely used Weather [43] benchmark. While the original dataset assumes uniform sampling periods across all channels, this assumption rarely holds in real-world environmental monitoring systems. In Weather-practical, we reorganize the channels to reflect heterogeneous sampling frequencies, based on the intrinsic temporal dynamics of each channel. For instance, fast-changing environmental channels such as wind velocity, precipitation, and solar radiation are sampled at 10-minute intervals, whereas more stable channels such as air pressure or vapor pressure are sampled at a coarser hourly rate. Mid-range channels—such as air temperature and specific humidity—are sampled at 20-minute or 30-minute intervals in accordance with their moderate temporal variability. This configuration, summarized in Table 7, mimics the asynchronous and heterogeneous acquisition patterns encountered in operational weather stations and provides a more realistic testbed for evaluating forecasting models under non-uniform input conditions.
@@
Monash-SolarWind-practical is derived from the Monash Forecasting Repository [10] and built upon high-frequency power generation data originally collected at 4-second intervals. In real-world energy monitoring systems, distinct sensors are used for solar and wind power, each exhibiting unique temporal behaviors and operational constraints. To replicate these real-world dynamics, we apply channel-specific downsampling: solar-related channels are resampled at 20-minute intervals, capturing smooth and gradual irradiance trends, while wind-related channels are resampled at 5-minute intervals to reflect the rapid fluctuations characteristic of wind dynamics. This restructured dataset captures modality-specific asynchrony and sampling heterogeneity, offering a challenging benchmark for assessing model robustness under realistic conditions where different signal types operate on distinct temporal resolutions.
@@
CHS is a real-world dataset collected from an LNG (Liquefied Natural Gas) carrier during ballast voyage. The dataset contains time series measurements from a total of 52 channels, each sampled at distinct frequencies based on their physical characteristics and operational requirements. Specifically, weather and navigation channels, such as latitude and wave period, are sourced from external systems and sampled at a coarse hourly rate. In contrast, onboard sensor channels are sampled more frequently to support real-time monitoring and control. These include machinery-related channels (e.g., Low Duty Compressor (LDC) suction pressure, main engine inlet pressure, and LNG consumption), sampled at 10-minute intervals, and cargo tank sensor channels (e.g., tank temperature at top, mid, and bottom levels, and tank pressure), sampled at 1-minute intervals. For experimental evaluation, we selected 10 key channels from the full set of 52, focusing on those most relevant to navigation and weather conditions, cargo tank thermodynamics, and machinery operations.
The LNG carrier is equipped with four independent cargo tanks. During voyage, ambient heat ingress causes the LNG stored in these tanks to partially vaporize, producing what is known as Boil-Off Gas (BOG). Efficient BOG management is critical: the gas can be used as fuel, reliquefied, or combusted via a Gas Combustion Unit (GCU), which converts it into CO2 before discharge. The operational objective is to minimize unnecessary combustion while maximizing energy utilization as fuel, all while maintaining tank pressure within safe operational bounds. The CHS dataset captures the dynamic interplay among external conditions, cargo tank behavior, and onboard machinery responses. It provides a foundation for predictive modeling of key control channels, including tank gas pressure and LDC suction pressure. Accurate prediction of these channels is important for safe and energy-efficient voyage execution under the channel-wise asynchronous sampling conditions found in real-world LNG cargo operations.
Figure 5: Overview of the LNG Cargo Handling System (CHS).
@@ A.2 Implementation Details
All experiments were conducted using a single NVIDIA RTX 3090 GPU with 24GB memory. We used the Adam optimizer across all models, with initial learning rates per dataset for fair comparison: $10^{-4}$ for ETT1 and SolarWind, $10^{-2}$ for Weather, and $10^{-3}$ for the CHS dataset. Training was run for a maximum of 10 epochs with early stopping based on validation performance. For models with a stackable encoder block, the number of encoder layers was set to 2. For Transformer-based models, the model dimension $d_{\text{model}}$ was searched over $\{128,256,512\}$, and the feedforward expansion ratio $d_{\text{ff}}$ / $d_{\text{model}}$ was selected from $\{1,2,4\}$. We adopted most baseline implementations from the TimesNet [44] repository. CrossGNN [15], originally provided in a separate repository, was reimplemented and integrated into the TimesNet framework for consistency. In our proposed ChannelTokenFormer (CTF), the number of channel tokens was selected from $\{1,2,3\}$, with the optimal value determined separately for each dataset.
@@
For baseline models, both training and testing were conducted on linearly interpolated data due to architectural constraints. To ensure a fair and consistent comparison, we preserved the original design of the baselines and trained them on interpolated inputs with MSE loss. For additional comparison, we also evaluated baseline models under the practical setting using forward-filled inputs. Empirically, we found that their performance remained largely consistent with that obtained using linear interpolation, suggesting that the interpolation method had limited influence on overall accuracy for these models (See Table 12).
In contrast, our proposed model, ChannelTokenFormer (CTF), is explicitly designed to operate on non-interpolated inputs that reflect the original, variable sampling periods of each channel. To ensure uniform input shape compatible with existing architectures, we apply forward-fill interpolation during dataset preprocessing; however, this step serves merely as a structural tool to meet input dimensionality requirements, not as a modeling assumption. During the channel-wise patch embedding phase, all forward-filled values are explicitly excluded: only the valid indices $L_{i}$, corresponding to truly observed data points, are used for computation.
To ensure fair evaluation, both the baselines and our model compute channel-aggregated metrics—CMSE and CMAE, as defined in Section 3—only over the valid indices $H_{i}$, which correspond to actually observed target values. Thus, although the input preprocessing methods differ, all models are evaluated on the same set of ground truth targets $(y_{1},\dots,y_{H_{i}})$. Note that under the practical setting, the number of valid test samples is inherently reduced, since the test set is subsampled according to the maximum sampling factor across channels. In summary, each model is trained under settings aligned with its architectural requirements, but all are evaluated uniformly based on their accuracy in predicting truly observed values.
@@
As described in Section 5, we set the input and prediction lengths for each dataset to reflect the temporal characteristics of the underlying signals. Instead of fixing the number of time steps across all channels, we define these lengths over a consistent time duration to accommodate channel-wise variations in both sampling period and dominant frequency components. For implementation and reporting convenience, however, we represent input and output lengths in terms of the number of steps corresponding to the channel with the smallest sampling period. The input and prediction lengths (in time steps, relative to the finest resolution channel) for each dataset are as follows: ETT1: 192 (input) $\rightarrow$ [192, 336, 768, 1152] (prediction), SolarWind: 576 $\rightarrow$ [288, 576, 864, 1152], Weather: 144 $\rightarrow$ [144, 288, 576, 864], CHS: 120 $\rightarrow$ [120, 240, 480, 840].
Example of Channel-wise Dynamic Patching
In our model, the number of local tokens per channel adapts to the characteristics of its signal: channels with finer temporal resolution or shorter dominant periods, as determined via FFT, result in more tokens, whereas coarser or slower signals result in fewer tokens within the same time window. For example, in the SolarWind dataset, wind power is recorded every 5 minutes and solar power every 20 minutes. Over a 2-day period, this corresponds to 576 time steps for wind power (dominant period: 48 steps) and 144 for solar power (dominant period: 18 steps). Assuming patching is performed based on each channel’s dominant period, this yields 576 / 48 = 12 local tokens for wind power and 144 / 18 = 8 for solar power.
@@
To evaluate robustness under structured missing conditions, we introduce block-wise missing intervals in the test-time inputs of two datasets. In the SolarWind dataset, each channel is assigned randomly positioned missing blocks, with each block spanning exactly one patch length. This ensures that at least one patch per channel is entirely unobserved, enabling a direct assessment of our model’s patch-level masking strategy at test time. During training, our model applies random patch masking, as shown in Table 15, training with a 0.4 masking ratio consistently improved model robustness across different missing ratios. In the ETT1 dataset, we insert random zero-filled missing blocks of varying lengths (5–20 time steps) into each channel. These settings are designed to test model robustness to short-range missing intervals. Detailed results under this setting are reported in Table 16.
@@ B.1 Background
In practical time series forecasting scenarios, multivariate signals often feature channel-wise asynchronous sampling periods across channels, stemming from both inter-system and intra-system heterogeneity in sensor design and operation. However, most existing forecasting models assume a regular and complete time grid, necessitating the use of interpolation to align all channels to a common temporal resolution prior to model ingestion. While this preprocessing step ensures input compatibility, it inevitably introduces spectral distortion [8, 9] by artificially synthesizing intermediate values, thereby altering the original frequency characteristics of the signals. Interpolation imposes artificial continuity by estimating unobserved values from nearby observations, which can fundamentally reshape both the temporal dynamics and spectral content of the signal. This appendix provides a theoretical explanation of how such interpolation-induced distortion arises in the frequency domain, as observed through the Fast Fourier Transform (FFT).
@@ B.3 Summary and Implication
Appendix C Ablation Study
@@ C.1 Additional Ablations
To complement the main experiments, we provide additional ablation studies highlighting the contribution of key architectural components in ChannelTokenFormer (CTF). As shown in Table 8, incorporating learnable channel-specific embeddings consistently improves performance across all datasets. This suggests that explicitly encoding channel identity through channel embeddings provides a strong inductive bias, particularly in heterogeneous multivariate settings.
@@
Setting / Datasets ETT1 SolarWind Weather CHS
CTF with channel embedding 0.398 0.400 0.269 0.282
CTF w/o channel embedding 0.404 0.399 0.283 0.285
@@
w/o Channel Dependence 0.412 0.464 0.435 0.484 0.457 0.507 0.482 0.532
w/o Dynamic patching 0.427 0.484 0.430 0.492 0.438 0.511 0.485 0.548
w/o Patch masking 0.435 0.486 0.486 0.525 0.536 0.564 0.574 0.593

### v2

@@ Appendix A Implementation Details
A.1 Dataset Configuration under Multi-Source Asynchrony
We conduct long-term forecasting experiments on four modified benchmark datasets adapted to our practical setting, as well as two real-world datasets. In this setting, each channel provides a different number of observations over the same temporal window, reflecting channel-specific sampling periods within the multivariate time series. This configuration more closely mirrors real-world conditions, where sensor deployment strategies and data acquisition policies vary across channels.
ETT1-practical & ETT2-practical (ETT1 & ETT2)
ETT1-practical and ETT2-practical are variants of the commonly used ETTm1 (Zhou et al., 2021) and ETTm2 datasets. Unlike their original versions, where all channels are sampled uniformly at 15-minute intervals, ETT1-practical and ETT2-practical introduce channel-specific sampling periods to reflect industrial environments, where sensors operate at different acquisition frequencies depending on their physical properties and monitoring requirements. Specifically, the channels are grouped into two categories: load-related channels and the oil-temperature (OT) channel. The load-related channels are downsampled to an hourly rate, representing slowly varying electrical load measurements, while the OT channel retains its original 15-minute resolution to capture high-frequency thermal dynamics.
@@ A.1 Dataset Configuration under Multi-Source Asynchrony
Weather-practical is constructed from the widely used Weather (Wu et al., 2021) benchmark. While the original dataset assumes uniform sampling periods across all channels, this assumption rarely holds in real-world environmental monitoring systems. In Weather-practical, we reorganize the channels to reflect heterogeneous sampling frequencies, based on the intrinsic temporal dynamics of each channel. For instance, fast-changing environmental channels such as wind velocity, precipitation, and solar radiation are sampled at 10-minute intervals, whereas more stable channels such as air pressure or vapor pressure are sampled at a coarser hourly rate. Mid-range channels—such as air temperature and specific humidity—are sampled at 20-minute or 30-minute intervals in accordance with their moderate temporal variability. This configuration, summarized in Table 7, mimics the asynchronous and heterogeneous acquisition patterns encountered in operational weather stations and provides a more realistic testbed for evaluating forecasting models under non-uniform input conditions.
@@
Monash-SolarWind-practical is derived from Australian Energy Market Operator (AEMO)  [^1] http://www.nemweb.com.au/ and the Monash Forecasting Repository (Godahewa et al., 2021) and built upon high-frequency power generation data originally collected at 4-second intervals. In real-world energy monitoring systems, distinct sensors are used for solar and wind power, each exhibiting unique temporal behaviors and operational constraints. To replicate these real-world dynamics, we apply channel-specific downsampling: solar-related channels are resampled at 20-minute intervals, capturing smooth and gradual irradiance trends, while wind-related channels are resampled at 5-minute intervals to reflect the rapid fluctuations characteristic of wind dynamics. This restructured dataset captures modality-specific asynchrony and sampling heterogeneity, offering a challenging benchmark for assessing model robustness under realistic conditions where different signal types operate on distinct temporal resolutions.
EPA-Air (EPA)
EPA-Air is sourced from air quality monitoring data (Chang et al., 2025) of the U.S. Environmental Protection Agency [^2] https://www.epa.gov/data, collects measurements from several monitoring regions across the United States. In real-world air quality monitoring, different channels are reported at heterogeneous temporal resolutions due to channel-specific measurement practices. To reflect these conditions, we assign channel-specific sampling periods: temperature-related variables are recorded at 1-hour intervals, fine particulate matter (PM2.5) is reported at 8-hour intervals, air quality indices (AQI) are provided at daily intervals, and ozone measurements are reported at weekly intervals. For our experiments, we selected four representative cities, Maricopa, Richmond, Los Angeles, and Hillsborough, from the original dataset that contains measurements from eight metropolitan areas. This real-world dataset preserves the inherent asynchrony and heterogeneity of environmental sensing, yielding a benchmark that mirrors practical conditions where meteorological variables, pollutant measurements, and derived indices are recorded on distinct temporal resolutions.
@@
CHS is a real-world dataset collected from an LNG (Liquefied Natural Gas) carrier during a ballast voyage. The dataset contains time series measurements from a total of 52 channels, each sampled at distinct frequencies based on their physical characteristics and operational requirements. Specifically, weather and navigation channels, such as latitude and wave period, are sourced from external systems and sampled at a coarse hourly rate. In contrast, onboard sensor channels are sampled more frequently to support real-time monitoring and control. These include machinery-related channels (e.g., Low Duty Compressor (LDC) suction pressure, main engine inlet pressure, and LNG consumption), sampled at 10-minute intervals, and cargo tank sensor channels (e.g., tank temperature at top, mid, and bottom levels, and tank pressure), sampled at 1-minute intervals. For experimental evaluation, we selected 10 key channels from the full set of 52, focusing on those most relevant to navigation and weather conditions, cargo tank thermodynamics, and machinery operations.
The LNG carrier is equipped with four independent cargo tanks. During voyage, ambient heat ingress into the cargo tanks causes the LNG stored in these tanks to partially vaporize, producing what is known as Boil-Off Gas (BOG). Efficient BOG management is critical: the gas can be used as fuel, reliquefied, or combusted via a Gas Combustion Unit (GCU), which converts it into CO2 before discharge. The operational objective is to minimize unnecessary combustion while maximizing energy utilization as fuel, all while maintaining tank pressure within safe operational bounds. The CHS dataset captures the dynamic interplay among external conditions, cargo tank behavior, and onboard machinery responses. It provides a foundation for predictive modeling of key control channels, including tank gas pressure and LDC suction pressure. Accurate prediction of these channels is important for safe and energy-efficient voyage execution under the channel-wise asynchronous sampling conditions found in real-world LNG cargo operations.
@@ A.2 Implementation Details
All experiments were conducted using a single NVIDIA RTX 3090 GPU with 24GB memory. We used the Adam optimizer across all models, with initial learning rates per dataset for fair comparison: $10^{-4}$ for ETT1, ETT2, SolarWind and EPA, $10^{-2}$ for Weather, and $10^{-3}$ for the CHS dataset. Training was run for a maximum of 10 epochs with early stopping based on validation performance. For models with a stackable encoder block, the number of encoder layers was set to 2. For Transformer-based models, the model dimension $d_{\text{model}}$ was searched over $\{128,256,512\}$, and the feedforward expansion ratio $d_{\text{ff}}$ / $d_{\text{model}}$ was selected from $\{1,2,4\}$. We adopted most baseline implementations from the TimesNet (Wu et al., 2023) repository. CrossGNN (Huang et al., 2023), t-PatchGNN (Zhang et al., 2024), Hi-Patch (Luo et al., 2025), BiTGraph (Chen et al., 2024b), originally provided in a separate repository, were reimplemented and integrated into the TimesNet framework for consistency. The hyperparameters of these baselines were aligned with those specified in their original repositories. Since ContiFormer (Chen et al., 2023) is an ODE-based model with prohibitively long training times, we reimplemented it within the TimesNet framework by approximating its ODE solver, following the structure of TimeXer (Wang et al., 2024e), to ensure both efficiency and consistency. In our proposed ChannelTokenFormer (CTF), the number of channel tokens was selected from $\{1,2,3\}$, with the optimal value determined separately for each dataset.
@@
For regular time series baseline models, both training and testing were conducted on linearly interpolated data due to architectural constraints. To ensure a fair and consistent comparison, we preserved the original design of the baselines and trained them on interpolated inputs with MSE loss. In contrast, our proposed model, ChannelTokenFormer (CTF), is explicitly designed to operate on non-interpolated inputs that reflect the original, variable sampling periods of each channel. To ensure uniform input shape compatible with existing architectures, we apply forward-fill interpolation during dataset preprocessing; however, this step serves merely as a structural tool to meet input dimensionality requirements, not as a modeling assumption. During the channel-wise patch embedding phase, all forward-filled values are explicitly excluded: only the valid indices $L_{i}$, corresponding to truly observed data points, are used for computation. For irregular and missing handling time series baselines, we followed their original frameworks by providing an input mask, which was used for both data processing and MSE computation. This procedure is exactly consistent with the forward-fill interpolation–based approach. However, since ContiFormer does not conform to other irregular frameworks and was reimplemented for execution, it follows the same procedure as the regular baselines. To ensure fair evaluation, both the baselines and our model compute channel-aggregated metrics—CMSE and CMAE, as defined in Section 3—only over the valid indices $H_{i}$, which correspond to actually observed target values. Thus, although the input preprocessing methods differ, all models are evaluated on the same set of ground truth targets $(y_{1},\dots,y_{H_{i}})$. Note that under the practical setting, the number of valid test samples is inherently reduced, since the test set is subsampled according to the maximum sampling factor across channels. In summary, each model is trained under settings aligned with its architectural requirements, but all are evaluated uniformly based on their accuracy in predicting truly observed values.
@@
As described in Section 5, we set the input and prediction lengths for each dataset to reflect the temporal characteristics of the underlying signals. Instead of fixing the number of time steps across all channels, we define these lengths over a consistent time duration to accommodate channel-wise variations in both sampling period and dominant frequency components. For implementation and reporting convenience, however, we represent input and output lengths in terms of the number of steps corresponding to the channel with the smallest sampling period. The input and prediction lengths (in time steps, relative to the finest resolution channel) for each dataset are as follows: ETT1, ETT2: 192 (input) $\rightarrow$ [192, 336, 768, 1152] (prediction), SolarWind: 576 $\rightarrow$ [288, 576, 864, 1152], Weather: 144 $\rightarrow$ [144, 288, 576, 864], EPA: 96 $\rightarrow$ [96, 192, 288, 384], CHS: 120 $\rightarrow$ [120, 240, 480, 840].
Details of Channel-wise Frequency-based Dynamic Patching
In dominant frequency detection via Fast Fourier Transform (FFT), each input channel is represented as a one-dimensional sequence $\mathbf{x}^{(i)}$, and its amplitude spectrum $\mathbf{A}^{(i)}=|\hat{\mathbf{x}}^{(i)}|$ is computed. Instead of relying on a predefined set of candidate periods, the dominant period is searched within a reasonable range (empirically up to $\sim$ 200), which covers common real-world cycles such as hourly, daily, and weekly periodicities. The frequency index $k^{*}_{i}$ with the highest amplitude in this range is identified and converted into the corresponding dominant period $p^{*}_{i}=\tfrac{L_{i}}{k^{*}_{i}}$. If the amplitude $\mathbf{A}^{(i)}$ at $k^{*}_{i}$ exceeds a fixed threshold (empirically set to the 70th percentile of $\mathbf{A}^{(i)}$), $p^{*}_{i}$ is regarded as a strong dominant period. In patch length assignment with relative sampling periods, the relative sampling period of channel $i$ is defined as $r_{i}=\tfrac{s_{i}}{s_{min}}$, where $s_{i}$ is the actual sampling period and $s_{min}$ is the minimum across all channels. The patch length $l_{i}$, representing the number of points per patch, is then assigned as $l_{i}=\lfloor p^{*}_{i}/r_{i}\rfloor$, ensuring consistency with both the detected periodicity and the sampling resolution. In the fallback mechanism, if no dominant frequency passes the 70th percentile threshold, a sampling-aware patching rule is applied that directly determines patch lengths based on each channel’s sampling rate, without enforcing periodicity-based constraints. This approach is designed for cases where periodicity is too weak or noisy to guide patching, and it prioritizes maintaining consistency with the intrinsic resolution of each channel.
Example of Channel-wise Frequency-based Dynamic Patching
For example, in the SolarWind dataset, wind power is recorded every 5 minutes and solar power every 20 minutes. Over a 2-day period, this corresponds to 576 time steps for wind power (resulting patch length: 48 steps) and 144 for solar power (resulting patch length: 18 steps). Assuming patching is performed based on each channel’s dominant period, this yields 576 / 48 = 12 local tokens for wind power and 144 / 18 = 8 for solar power. Detailed calculations are provided below.
  • Relative sampling periods: $r_{1}=1,r_{2}=4$
  • FFT-detected dominant periods: $p^{*}_{1}=48,p^{*}_{2}=72$
  • Resulting patch lengths: $l_{1}=\lfloor p^{*}_{1}/r_{1}\rfloor=\lfloor 48/1\rfloor=48$ , $l_{2}=\lfloor p^{*}_{2}/r_{2}\rfloor=\lfloor 72/4\rfloor=18$
  • Number of patches:
    • $P_{ch1}=\lfloor L/l_{1}\rfloor=\lfloor 576/48\rfloor=12$
    • $P_{ch2}=\lfloor L/(l_{2}\cdot r_{2})\rfloor=\lfloor 576/(18\cdot 4)\rfloor=8$
@@
To evaluate robustness under structured missing conditions, we introduce block-wise missing intervals in the test-time inputs of three datasets, including real-world EPA air monitoring dataset. In the SolarWind dataset, each channel is assigned randomly positioned missing blocks, with each block spanning exactly one patch length. This ensures that at least one patch per channel is entirely unobserved, enabling a direct assessment of our model’s patch-level masking strategy at test time. During training, our model applies random patch masking, as shown in Table 18, training with a 0.4 random masking ratio consistently improved model robustness across different missing ratios. In the ETT1 and EPA datasets, we insert random zero-filled missing blocks of smaller varying lengths (5 to 20 time steps) into each channel. These settings are designed to test model robustness to short-range missing intervals. For the ETT1 missing setting, we compared against regular time series baselines, which are better suited to ETT data. For the EPA missing setting, we selected representative baselines, Hi-Patch and ContiFormer. BiTGraph was excluded from the EPA experiments as it failed to converge properly. Also, we selected LA and Hillsborough to evaluate in a more challenging environment. Detailed results under this setting are reported in Table 19 and Table 20.
@@ B.1 Background
In practical time series forecasting scenarios, multivariate signals often feature channel-wise asynchronous sampling periods across channels, stemming from both inter-system and intra-system heterogeneity in sensor design and operation. However, most existing forecasting models assume a regular and complete time grid, necessitating the use of interpolation to align all channels to a common temporal resolution prior to model ingestion. While this preprocessing step ensures input compatibility, it inevitably introduces spectral distortion (Feng et al., 2004; Fitzgerald & Anderson, 1992) by artificially synthesizing intermediate values, thereby altering the original frequency characteristics of the signals. Interpolation imposes artificial continuity by estimating unobserved values from nearby observations, which can fundamentally reshape both the temporal dynamics and spectral content of the signal. This appendix provides a theoretical explanation of how such interpolation-induced distortion arises in the frequency domain, as observed through the Fast Fourier Transform (FFT).
@@ B.2 Linear Interpolation-induced Frequency Domain Distortion
B.3 Summary and Implication
Appendix C Further Analysis
@@ C.1 Additional Ablations
To complement the main experiments, we provide additional ablation studies highlighting the contribution of key architectural components in ChannelTokenFormer (CTF). As shown in Table 8, incorporating learnable channel-specific embeddings consistently improves performance across all datasets. This suggests that explicitly encoding channel identity through channel embeddings provides a strong inductive bias, particularly in heterogeneous multivariate settings. Furthermore, as shown in Table 9, setting patch lengths to match each channel’s periodic characteristics, rather than using a fixed patch length, has a clear impact on forecasting performance. This underscores the importance of reflecting channel-specific sampling periodicities when handling asynchronous time series.
@@
Setting / Dataset ETT1 ETT2 SolarWind Weather EPA CHS
CTF with channel embedding 0.398 0.377 0.400 0.269 0.775 0.282
CTF w/o channel embedding 0.404 0.379 0.399 0.283 0.781 0.285
Table 9: Impact of patch length on forecasting performance. Results are reported in CMSE, averaged over all prediction lengths. In general, frequency-guided patching improves performance on most datasets, suggesting that adapting patch lengths to each channel’s periodic structure is beneficial.
Patch length / Dataset ETT1 ETT2 SolarWind Weather EPA CHS
FFT-based 0.398 0.377 0.400 0.274 0.775 0.282
8 0.404 0.378 0.405 0.289 0.817 0.300
16 0.401 0.378 0.400 0.289 0.807 0.301
32 0.424 0.383 0.399 0.306 0.799 0.307
@@
w/o Channel Dependence 0.408 0.463 0.429 0.484 0.474 0.521 0.522 0.551
w/o Dynamic patching 0.452 0.496 0.474 0.515 0.494 0.536 0.517 0.562
w/o Patch masking 0.412 0.465 0.434 0.484 0.458 0.508 0.482 0.533